### v1

# Video-3D LLM: Learning Position-Aware Video Representation for 3D Scene Understanding

Abstract
The rapid advancement of Multimodal Large Language Models (MLLMs) has significantly impacted various multimodal tasks. However, these models face challenges in tasks that require spatial understanding within 3D environments. Efforts to enhance MLLMs, such as incorporating point cloud features, have been made, yet a considerable gap remains between the models’ learned representations and the inherent complexity of 3D scenes. This discrepancy largely stems from the training of MLLMs on predominantly 2D data, which restricts their effectiveness in comprehending 3D spaces. To address this issue, in this paper, we propose a novel generalist model, i.e., Video-3D LLM, for 3D scene understanding. By treating 3D scenes as dynamic videos and incorporating 3D position encoding into these representations, our Video-3D LLM aligns video representations with real-world spatial contexts more accurately. Additionally, we have implemented a maximum coverage sampling technique to optimize the balance between computational costs and performance efficiency. Extensive experiments demonstrate that our model achieves state-of-the-art performance on several 3D scene understanding benchmarks, including ScanRefer, Multi3DRefer, Scan2Cap, ScanQA, and SQA3D. Our code is available at https://github.com/LaVi-Lab/Video-3D-LLM.
1 Introduction
The rapid development of Multimodal Large Language Models (MLLMs) [29, 38, 34, 28, 43, 15, 2, 18, 35, 53] has demonstrated substantial capabilities in various multi-modal tasks, attracting significant attention from both academia and industry sectors. However, despite these advancements, recent studies [33, 37, 32, 5] indicate that current MLLMs face challenges when addressing tasks that necessitate spatial understanding and reasoning in 3D environments.
Figure 1: Comparison of previous work and our method: (a) Previous 3D LLMs are initialized on MLLMs trained solely on image-text pairs, and learn point cloud or voxel representations via fine-tuning on 3D scenes. The 3D point clouds are reconstructed from RGB-D videos. (b) Our method directly utilizes video frames and 3D coordinates as input, where the 3D coordinates are converted from depths through coordinate transformation. We then transfer the ability of video understanding to 3D scene understanding by injecting position information into video representations.
Recent studies [21, 11, 47, 14, 57, 22, 45, 19, 25] have focused on adapting MLLMs for enhanced 3D scene understanding. As depicted in Figure 1 (a), these approaches develop comprehensive 3D scene-level representations using a variety of techniques. These include harnessing features from point clouds [47, 14, 11], converting attributes from multi-view images into 3D formats [21, 57, 19], and exploiting characteristics from recognized objects [23, 45, 22, 25].
Although significant advancements have been achieved, a noticeable gap exists between the representations learned by MLLMs and the complexity of 3D scenes. This gap stems from the fact that MLLMs are primarily trained on fundamentally different data types, namely 2D images. While it is possible to further finetune MLLMs with 3D data, such as point clouds or voxels, the limited availability of labeled 3D scene data poses a challenge. Consequently, the 2D visual knowledge embedded in MLLMs offers limited assistance in understanding 3D environments.
In parallel, the abundance of video data has spurred interest in adapting Video LLMs to different domains, e.g., 3D question answering [35, 28, 33] and robotic manipulation [60, 27, 46]. These methods benefit from extensive internet video datasets and pre-trained video models, revealing the immense potential for extending video modality to 3D modeling. However, as early attempts, they are still far from creating a model capable of handling diverse 3D tasks. Moreover, the absence of integrated spatial information—such as 3D locations and spatial relationships—in video representations constrains their capability to fully comprehend the 3D physical world. For instance, tasks that require an intricate understanding of 3D spatial relationships cannot rely solely on RGB data. This limitation underscores the necessity for incorporating more comprehensive spatial modeling into Video LLMs to enhance their effectiveness in 3D applications.
Therefore, in this paper, we propose a generalist model for 3D scene understanding, namely Video-3D LLM. As shown in Figure 1, our model is based on a Video LLM framework and processes 3D videos, i.e., video frames accompanied by the corresponding 3D spatial coordinates. Typically, as the current 3D scenes are mostly reconstructed from RGB-D videos, we can seamlessly project the depth images into the global coordinate system. To establish the correspondence in visual appearance and position information, we learn position-aware video representations by injecting 3D coordinates into video features. Subsequently, we encode the coordinate to 3D position encoding and add it to the video representations, serving as the input for the Video LLM.
Our model offers several significant advantages. Firstly, it aligns video representations with their real-world spatial contexts, thereby equipping our model to handle various 3D tasks such as 3D visual grounding, 3D dense captioning, and 3D question answering. Secondly, it maintains both temporal and spatial contextual information in the video data, which helps to reduce the discrepancy between the pre-training data and actual 3D scenes. Additionally, we have developed a maximum coverage sampling strategy for frame selection. This approach views frame selection as a maximum coverage problem and adopts a greedy algorithm for its resolution. This strategy ensures the selection of the most informative frames, thus improving the model’s capacity to discern diverse and essential spatio-temporal features within the video, while also ensuring efficient inference performance.
Our approach is to train a single model in the multi-task manner on varying 3D scene understanding tasks, including 3D question answering, 3D dense captioning, and 3D visual grounding. Extensive experiments demonstrate that our Video-3D LLM achieves state-of-the-art performance on five 3D scene understanding benchmarks, i.e., ScanRefer [6], Multi3DRefer [52], Scan2Cap [7], ScanQA [3] and SQA3D [36]. Notably, our method surpasses the previous state-of-the-art LLaVA-3D [57] by using only 26% of its 3D data (223k vs. 859k), achieving improvements of 4.1% Acc@0.25 on ScanRefer, 4.6 CIDEr@0.5IoU on Scan2Cap, 2.9% EM on ScanQA, and 3.0% EM on SQA3D. The impressive performance reveals the immense potential for adapting video models to 3D modality, establishing a new paradigm in 3D scene understanding.
2 Related Work
3D Scene Understanding. The growing interest in 3D scene understanding has spurred the development of various tasks, such as: (1) 3D visual grounding [1, 6, 52, 56, 49, 54, 9, 58, 51, 24, 44], which involves pinpointing a target object within a 3D environment based on a textual description. (2) 3D dense captioning [7, 10, 12], which requires the model to localize and provide detailed captions for objects within a 3D scene. (3) 3D question answering [3, 36], which focuses on answering questions related to 3D scenes. These 3D vision-language tasks differ from 2D tasks [13, 20] as they necessitate enhanced modeling of spatial positions and relative relationships. To meet these challenges, various approaches have been introduced, addressing aspects such as view changes [24], spatial modeling [54, 9], and fine-grained alignment [52, 58, 51, 10, 12]. However, these methods are typically specialized for specific tasks and lack generalizability. In contrast, our work aims to build a generalist model capable of handling diverse 3D-VL tasks, thereby offering a more flexible and widely applicable solution in the field.
LLMs for 3D Scene Understanding. Recently, there has been a growing interest in integrating 3D information in LLMs [21, 14, 45, 23, 47, 22, 57, 11, 25], which advances the progress of 3D scene understanding. 3D-LLM [21] introduces the LLM-based model for 3D physical world, which takes 3D features from rendered 2D images as input. PointLLM [47] utilizes a point encoder with a strong LLM for point cloud understanding. LL3DA [11] leverages a Q-former to extract useful information from point cloud features, endowing humans with the capability to interact in 3D environments. Grounded 3D-LLM [14] further introduces a projection module based on 3D detectors, which allows for generating object proposals from point-level features. Chat3D[45], LEO [23], and ChatScene [22] take use of off-the-shelf 3D detectors for proposal generation, and then incorporate the object-centric representations into LLMs. LLaVA-3D [57] introduces 3D-patch representations, which aggregate 2D-patch features in voxel space. Robin3D [25] tries to enhance 3D scene understanding via data generation. It is important to note that existing 3D Large Language Models (3D LLMs) typically transform 3D scenes into voxel-level or point cloud-level 3D representations as input for modeling purposes. However, these approaches create a disconnect with pre-trained multi-modal large language models (MLLMs), which are primarily trained on extensive 2D datasets, such as images, and only fine-tuned on a limited amount of 3D scene data. To address this challenge, our method incorporates 3D information (e.g., coordinates shown in Figure 1 (b)) into a new video representation. This enhancement maximizes the use of pre-trained 2D Video LLMs, leveraging their full potential.
Video-Language Models for 3D Understanding. We have witnessed the rapid development of Video LLMs [50, 53, 35, 30, 31]. There is also a trend of leveraging Video LLMs for 3D tasks, including 3D question answering [35, 28, 33] and robotic manipulation [60, 27, 46]. LLaVA-OneVision [28] and Oryx MLLM [35] incorporate 3D question-answering datasets into instruction data, which deliver competent results on 3D question-answering tasks. However, these models do not capture detailed 3D spatial information, which limits their performance in addressing other 3D tasks that require precise spatial alignments. Furthermore, recent work [33] has emphasized the importance of identifying key object correspondences across frames through visual prompting. In contrast, our approach directly incorporates 3D positional information into video representation learning, which enhances the capability to tackle more complex tasks that demand a thorough spatial understanding of the 3D environment.
Figure 2: The overview of the model architecture. (a) shows the integration of video sequence and global coordinates for creating position-aware video representations. (b) and (c) detail the examples of 3D dense captioning and 3D visual grounding, respectively. Our approach can generalize well to other 3D tasks.
3 Method
We propose a generalist model for 3D scene understanding, namely Video-3D LLM, built upon a pre-trained Video LLM [53]. Our model includes a visual encoder, a 3D position encoding module, and a Video LLM backbone. To effectively utilize video representation for 3D scene understanding, it is necessary to first convert 3D scenes into sequences of frames. Given that entire frame sequences can be redundant, implementing an effective frame selection strategy is crucial, as it significantly influences both performance and computational efficiency. Subsequently, our goal is to enhance the Video LLM with position awareness. This is achieved by encoding spatial coordinates into the 3D position encoding (3D-PE) and integrating them into the video representation learning. In this section, we detail three key components of our approach: the frame sampling strategy (3.1), position-aware video representation (3.2), and the training objective (3.3).
3.1 Frame Sampling Strategy
Converting 3D scenes into video sequences presents two primary challenges: (1) Due to GPU memory constraints, the Video LLM can only process a limited number of frames at a time. This necessitates the sampling of a subset of frames from the extensive raw video sequence to manage resources effectively. (2) It is crucial for the video sequence to encompass as much of the entire 3D scene as possible, since any omission of scene content could result in a significant and irreversible decline in model performance. To address these challenges, we introduce a maximum coverage strategy for frame sampling. This approach involves preprocessing the selected frames offline and applying the strategy consistently during both the training and inference phases to ensure comprehensive scene coverage and efficient memory usage.
Frame Sampling as Maximum Coverage Problem. Given a raw RGB-D video, each frame captures a portion of the 3D scene. We aim to select fewest possible frames that maximize coverage of the 3D scene, which could be formulated as a maximum coverage problem.
Formally, let $F=\{f_{1},f_{2},\ldots,f_{n}\}$ represent the set of all frames, and $V=\{v_{1},v_{2},\ldots,v_{m}\}$ represent the set of all voxels in the 3D scene. Each frame $f_{k}$ covers a subset of voxels $V_{k}\subseteq V$, which are discretized by the coordinates. The objective is to find a subset of frames $S\subseteq F$ such that the union of covered voxels $\bigcup_{f_{k}\in S}V_{k}$ is maximized.
Greedy Solution. Since the maximum coverage problem is NP-hard, we employ a greedy algorithm to solve it, which can obtain an approximation ratio of $1-1/e$ [26]. As illustrated in the Algorithm 1, the approach iteratively selects the frame with the largest increase in uncovered voxel coverage. Frames are added until the desired number is reached or the coverage ratio exceeds a predefined threshold. This stop condition ensures a balance between computational efficiency and the coverage for varying scenes.
Algorithm 1 Maximum Coverage Sampling
0:  Set of frames $F=\{f_{1},f_{2},\ldots,f_{n}\}$, voxel sets $V_{k}$ for each frame $f_{k}$, budget $K$
0:  Subset $S\subseteq F$ maximizing voxel coverage
1:  Initialize $S\leftarrow\emptyset$
2:  Initialize $U\leftarrow\emptyset$Set of covered voxels}
3:  while size of $S$ is less than $K$ do
4:     Select $f^{*}=\arg\max_{f_{k}\in F\setminus S}|V_{k}\setminus U|$
5:     Add $f^{*}$ to $S$
6:     Update $U\leftarrow U\cup V_{f^{*}}$
7:     if  Stop condition is met then
8:        $\mathbf{break}$
9:     end if
10:  end while
11:  return  $S$
3.2 Position-Aware Video Representation
After completing frame sampling, we obtain a sequence of RGB frames, depth images, and the camera’s intrinsic and extrinsic parameters. To create a position-aware video representation, we first transform the depth information into 3D coordinates within a global coordinate system. We then encode the visual embeddings along with the 3D position encodings (3D-PE) to enhance spatial awareness.
Camera Coordinate Transformation. Given a depth image $D\in\mathbb{R}^{H\times W}$, an extrinsic matrix $T_{\mathrm{camera}}^{\mathrm{global}}\in\mathbb{R}^{4\times 4}$, and a camera intrinsic matrix $K\in\mathbb{R}^{3\times 3}$, we can calculate the global coordinates $p_{\mathrm{global}}$ of pixel position $(i,j)$ using the following formula:
$\begin{bmatrix}p_{\mathrm{global}}\\ 1\end{bmatrix}=T_{\mathrm{camera}}^{\mathrm{global}}\cdot\begin{bmatrix}D(i,j)\cdot K^{-1}\cdot\begin{bmatrix}j\\ i\\ 1\end{bmatrix}\\ 1\end{bmatrix},$ (1)
as a result, we obtain a sequence of frames $\{f_{k}\}_{k=1}^{l}$ along with a set of coordinates in the global coordinate system $\{c_{k}\}_{k=1}^{l}$, where each $f_{k}$ and $c_{k}$ are in $\mathbb{R}^{H\times W\times 3}$.
Visual Embedding. We first encode each frame into visual embeddings via a Vision Transformer (ViT) [17]. In concrete, given a frame $f_{i}\in\mathbb{R}^{H\times W\times 3}$, the image will first be split into a series of patches at the patch size $P$, which are then fed into the ViT to produce visual embeddings $\mathbf{e}^{img}_{k}\in R^{H^{\prime}\times W^{\prime}\times d}$, where $H^{\prime}=\left\lfloor\frac{H}{P}\right\rfloor$, $W^{\prime}=\left\lfloor\frac{W}{P}\right\rfloor$ and $d$ is the feature dimension.
3D Position Encoding. Since the video frame is divided into image patches, we need to pool the coordinate information of each image patch. For each global coordinate map $c_{k}\in R^{H\times W\times 3}$, we divide the coordinates into patches identical to those of the image patches. Subsequently, we compute the average coordinate for each patch, which could be denoted as:
$c_{k}^{\prime}(i,j)=\frac{1}{P^{2}}\sum_{(u,v)\in\text{patch at }(i,j)}c_{k}(u,v),$ (2)
where each $c_{k}^{\prime}\in R^{H^{\prime}\times W^{\prime}\times d}$. Due to the small size of the patches, the averaged coordinates retain sufficiently precise positional information. We also explored alternative coordinate pooling methods in our ablation study.
We adopt sinusoidal position encoding [41] to encode the coordinates. For a 3D coordinate $(x,y,z)$, we first map the coordinate onto a discrete grid. The encoding for the $x$ coordinate is then defined as:
$\text{PE}(x,2i)$ $=\sin\left(\frac{x}{10000^{2i/\left\lfloor\frac{d}{3}\right\rfloor}}\right),$ (3)
$\text{PE}(x,2i+1)$ $=\cos\left(\frac{x}{10000^{2i/\left\lfloor\frac{d}{3}\right\rfloor}}\right).$ (4)
Similar calculations are applied to the $y$ and $z$ coordinates. The PE of $(x,y,z)$ are concatenated to obtain the final coordinate embeddings $\mathbf{e}^{\mathrm{coord}}_{k}\in R^{H^{\prime}\times W^{\prime}\times d}$. Lastly, the coordinate embeddings are added to the visual embeddings to form the position-aware video representations, denoted by:
$\mathbf{e}^{vis}_{k}=\mathbf{e}^{img}_{k}+\mathbf{e}^{coord}_{k}.$ (5)
3.3 Training Objective
Our approach is to build a generalist model which can handle multiple tasks with the single learned model. Our model is trained using a combined dataset that encompasses a variety of 3D scene understanding tasks in the multi-task manner. During training, we uniformly sample batches where each batch contains data from a single task type. For general 3D scene understanding tasks, such as 3D question answering and 3D dense captioning, we use cross-entropy loss to supervise text generation. For batches sampled from the 3D visual grounding task, to locate more accurately, we use the designed 3D visual grounding loss doing the multi-task training.
Cross-Entropy Loss. Given the position-aware video representation and a textual instruction, the language modeling objective aims to optimize the cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$:
$\mathcal{L}_{\mathrm{CE}}=-\sum\log(y|\{\mathbf{e}_{k}^{vis}\}_{k=1}^{l},\{\mathbf{e}_{k}^{text}\}_{k=1}^{q}),$ (6)
where $y$ is the ground truth response, $\{\mathbf{e}_{k}^{vis}\}_{k=1}^{l}$ are position-aware video representations for the video and $\{\mathbf{e}_{k}^{text}\}_{k=1}^{q}$ are text embeddings.
[…]
Method 3D Generalist ScanRefer Multi3DRef Scan2Cap ScanQA SQA3D
Acc@0.25
Acc@0.5
F1@0.25
F1@0.5
B-4@0.5
C@0.5
C EM EM
Expert Models
ScanRefer [6] 37.3 24.3
MVT [24] 40.8 33.3
3DVG-Trans [54] 45.9 34.5
ViL3DRel [9] 47.9 37.7
M3DRef-CLIP [52] 51.9 44.7 42.8 38.4
Scan2Cap [7] 22.4 35.2
ScanQA [3] 64.9 21.1 47.2
3D-VisTA [58] 50.6 45.8 34.0 66.9 69.6 22.4 48.5
2D LLMs
Oryx-34B [35] 72.3
LLaVA-Video-7B [53] 88.7 48.5
3D LLMs
3D-LLM(Flamingo)
[21]
21.2 59.2 20.4
3D-LLM(BLIP2-flant5)
[21]
30.3 69.4 20.5
Chat-3D [45] 53.2
Chat-3D v2  [22] 42.5 38.4 45.1 41.6 31.8 63.9 87.6 54.7
LL3DA [11] 36.0 62.9 76.8
SceneLLM [19] 80.0 27.2 53.6
LEO [23] 38.2 72.4 101.4 21.5 50.0
Grounded 3D-LLM [14] 47.9 44.1 45.2 40.6 70.6 35.5 72.7
PQ3D [59] 57.0 51.2 50.1 36.0 80.3 47.1
ChatScene [22] 55.5 50.2 57.1 52.4 36.3 77.1 87.7 21.6 54.6
LLaVA-3D [57] 54.1 42.4 41.1 79.2 91.7 27.0 55.6
Video-3D LLM (MC) 57.9 51.2 57.9 52.4 40.2 80.0 100.5 29.5 57.7
Video-3D LLM (Uniform) 58.1 51.7 58.0 52.7 41.3 83.8 102.1 30.1 58.6
Table 1: Overall performance comparison. “Expert models” are customized for specific tasks through task-oriented heads. “3D Generalist” means the model can perform multiple 3D tasks in a single model. LLaVA-Video is assessed in a zero-shot setting.
3D Visual Grounding Objective. Previous studies [21, 56] have demonstrated that directly outputting 3D bounding boxes is quite challenging for LLM. To enable our model to perform 3D visual grounding, we follow previous [22, 24, 58] work to model the task as a proposal classification task. As illustrated in Fig 2 (c), given a list of object proposals, we extract object features for each object from the visual embeddings. Specifically, for each object $b_{i}$, we check each patch to see if more than 50% of its points are contained within $b_{i}$, and then apply average pooling to the features of all selected patches, which can be formalized as:
$\mathbf{e}^{\text{obj-2d}}_{i}=\underset{{\text{if the percentage of points in}b_{i}>50\%}}{\text{AveragePooling}}(\mathbf{e}_{k}^{img}(u,v)).$ (7)
[…]
$\mathcal{L}_{Grd}=\frac{\exp(f(\mathbf{e}^{\text{obj}}_{k})\cdot g(h)/\tau)}{\sum_{i}\exp(f(\mathbf{e}^{\text{obj}}_{i})\cdot g(h)/\tau)},$ (8)
where $k$ is the index of the ground truth object, $f$ and $g$ are two-layer learnable MLPs, and $\tau$ is the temperature.
4 Experiments
In this section, we first compare the overall performance of Video-3D LLM with top-tier models and also investigate the effectiveness of all components.
4.1 Experimental Setup
Datasets. We conduct experiments across five 3D scene understanding benchmarks. For visual grounding, we test our model on ScanRefer [6] and Multi3DRefer [52], which require localizing objects in single-target and multiple-target scenarios, respectively. For dense captioning, we utilize the Scan2Cap [7] benchmark, which involves densely generating descriptions for all objects in 3D scenes. For question answering, we use the ScanQA [3] for spatial scene understanding and SQA3D [36] for situated reasoning. All these datasets are sourced from the ScanNet [16], a richly annotated RGB-D video dataset containing 1,513 scans in 3D scenes. We pre-process video frames for each scan at 3 FPS and extract the corresponding extrinsic and camera intrinsic parameters. For evaluation, we follow previous work [22, 57, 14] to adopt the validation sets for ScanRefer, Multi3DRefer, Scan2Cap, and ScanQA, and adopt the test set for SQA3D.
Metrics. We adopt widely used evaluation metrics for each of these benchmarks. For ScanRefer [6], we report thresholded accuracy metrics, specifically Acc@0.25 and Acc@0.5, where a prediction is considered correct if its Intersection over Union (IoU) with the ground truth exceeds 0.25 and 0.5, respectively. For Multi3DRefer [52], which involves grounding a variable number of target objects, we use the F1 score at IoU thresholds of 0.25 and 0.5. For Scan2Cap [7], we apply CIDEr@0.5IoU and BLEU-4@0.5IoU (denoted as C@0.5 and B-4@0.5), combining traditional image captioning metrics with IoU between predicted and reference bounding boxes. For ScanQA [3], we use CIDEr [42] and exact match accuracy, referred to as C and EM, respectively. Finally, for SQA3D [36], we evaluate performance using exact match accuracy (EM).
Implementation Details. We build Video-3D LLM based on the LLaVA-Video 7B [53], an open-sourced video LLM based on the QWen2-7B [48]. We use the Adam optimizer to train our model for one epoch with a batch size of 16 and a warmup ratio of 0.03. During the warmup phase, the learning rates peak at 1e-5 for the LLM and 2e-6 for the vision encoder. All experiments are conducted on 8 A100-80G GPUs. For 3D visual grounding and dense captioning, in training, we use the ground truth objects as the candidates. While in inference, we follow [22, 23] to employ Mask3D [39] to generate object proposals. The temperature $\tau$ for InfoNCE loss is 0.07.
4.2 Comparison with State-of-the-art Methods
4.2.1 Comparison Baselines
For a comprehensive comparison, we include both expert models designed for specific tasks and LLM-based models.
Expert Models. For ScanRefer [6], we compare our method with ScanRefer [6], MVT [24], 3DVG-Trans [54], ViL3DRel [9]. M3DRef-CLIP [52] further extends 3D grounding capabilities to multi-target scenarios. Scan2Cap [7] and ScanQA [3] provide initial benchmarks for the Scan2Cap and ScanQA datasets, respectively. 3D-VisTA [58] is pre-trained on large-scale scene-text pairs and then finetuned on specific tasks.
2D LLMs. Oryx [35] has included the ScanQA dataset in its training stage. In addition, we test the zero-shot performance of LLaVA-Video [53].
3D LLMs. 3D-LLM [21] is the first LLM-based model for 3D scene undersanding. SceneLLM and LL3DA [11] enrich the 3D representations with point cloud features. Chat3D [45], LEO [23], and ChatScene [22] incorporate object representations into 3D LLMs. Grounded 3D-LLM [14], PQ3D [59] and LLaVA-3D [57] deliver impressive results on 3D visual grounding by co-training a 3D detector. All these methods apply 3D point cloud features or projecting multi-view image features into 3D space, while ours directly works on the video representations.
4.2.2 Results
Frame Number Sampling Strategy Inference Time ScanRefer Multi3dRefer Scan2Cap ScanQA SQA3D
Acc@0.25 Acc@0.5 F1@0.25 F1@0.5 B-4@0.5 C@0.5 C EM EM
Fixed Frame Number
8 Uniform 309ms 48.93 43.50 49.80 45.40 37.34 68.82 94.98 27.57 56.77
MC 53.47 47.41 53.55 48.54 38.77 73.08 96.37 28.00 56.97
16 Uniform 537ms 55.42 49.17 54.95 49.82 39.39 76.96 99.86 28.96 57.70
MC 56.46 50.11 56.65 51.39 39.59 76.84 100.63 29.49 57.82
32 Uniform 1050ms 58.11 51.72 58.02 52.68 41.30 83.76 102.06 30.09 58.56
MC 58.27 51.68 57.93 52.50 40.32 81.58 102.33 30.35 59.25
Adaptive Frame Number
$\approx$18 MC 527ms 57.86 51.18 57.87 52.40 40.18 80.00 100.54 29.50 57.72
Previous SOTA
LLaVA-3D [57] 433ms 54.1 42.4 41.1 79.2 91.7 27.0 55.6
Table 2: Ablation study for the effect of frame sampling strategy. “MC” represents maximum coverage sampling. “MC” denotes sampling frames until over 95% of the scene’s voxels are covered or a maximum of 32 frames is reached.
We present the overall comparison with leading methods in Table 1. “Video-3D LLM (Uniform)” is trained using uniform sampling with 32 frames, while “Video-3D LLM (MC)” is trained using maximum coverage sampling with a coverage ratio of 95% and a maximum frame number of 32. “Video-3D LLM (Uniform)” achieves state-of-the-art performance on a variety of tasks including 3D visual grounding, 3D dense captioning, and 3D question answering, while “Video-3D LLM (MC)” delivers similar results with only half the inference time (527ms vs. 1050 ms).
3D Visual Grounding. For 3D visual grounding, we follow previous work [24, 58, 22] to detect all objects using Mask3D [39] and then make predictions over the object proposals. Our model achieves the highest accuracy, with Acc@0.25 at 58.1% and Acc@0.5 at 51.7% on ScanRefer, and F1@0.25 at 58.0% and F1@0.5 at 52.7% on Multi3DRefer. As previous 3D LLMs either use detected object proposals (e.g., ChatScene, Chat3D) or train an additional grounding module based on a 3D detector (e.g., Grounded 3d-LLM, PQ3D, LLaVA-3D), we can compare with these 3D LLMs fairly. Specifically, our model improves Acc@0.25 by 2.6% on ScanRefer and F1@0.25 by 0.9% on Multi3DRefer compared to ChatScene, which uses the same object proposal as ours.
3D Dense Captioning. Following the previous setting [23, 22], we generate captions for each detected object proposal. Our method demonstrates superior performance in Scan2Cap, achieving 41.3 at B-4@0.5 and 83.8 at C@0.5. The results reveal that our method connects video content with its position information by injecting 3D-PE into video representations.
3D Question Answering. Our model outperforms best competitors with 30.1% EM on ScanQA and 58.6% EM on SQA3D, which could be attributed to the strong representations inherited from Video LLMs. It is worth noting that existing Video LLMs achieve competent results in a zero-shot manner, suggesting that current 3D question answering tasks may not sufficiently address the challenges of spatial understanding in 3D scenes.
3D-PE Coord. ScanRefer Scan2Cap ScanQA
Acc@0.25 Acc@0.5 C@0.5 EM
None Avg 57.50 50.84 31.03 30.03
MLP 59.63 52.98 76.23 29.62
Sin 58.11 51.72 83.76 30.09
Sin Center 57.53 51.06 80.88 29.39
Min-Max 58.05 51.77 82.75 30.18
Avg 58.11 51.72 83.76 30.09
Table 3: Ablation study for the effect of coordinate encoding. “Coord.” means the method for aggregating the coordinates.
4.3 Ablation Study
Effectiveness of Frame Sampling. In Table 2, we assess the effectiveness of the frame sampling strategy across varying numbers of frames. To evaluate the model’s efficiency, we calculated the average inference speed on ScanQA. In the fixed frame number setting, we sample frames until reaching the desired number. As shown in the table, the model’s performance improves with an increasing number of frames, though inference time also rises. Video-3D LLM surpasses many previous methods with only 8 frames, demonstrating that our position-aware video paradigm can effectively model 3D scene understanding tasks. With 8 and 16 frames, maximum coverage sampling (“MC”) significantly enhances performance for all tasks by capturing a more complete 3D scene. Notably, with 8 frames, it improves Acc@0.25 by $4.54$ on ScanRefer and C@0.5 by $4.26$ on Scan2Cap compared to the uniform sampling strategy (“Uniform”). With 32 frames, the “Uniform” strategy shows comparable results with “MC”, as this number of frames is more than enough to cover the 3D scene. In the adaptive frame number setting, frames are sampled until over 95% of the scene’s voxels are covered or a maximum of 32 frames is reached. This allows for flexible adjustment based on scene size. As shown in the results, “MC” uses an average of $18$ frames across all scenes, achieving similar performance to the 32-frame uniform strategy while offering better inference speed. Meanwhile, compared to LLaVA-3D [57], “MC” shows superior performance while achieving similar inference speed.
Patch Size Loss ScanRefer Multi3DRefer
Acc@0.25 Acc@0.5 Acc@0.5 Acc@0.5
14 InfoNCE 56.44 50.08 56.31 51.05
27 InfoNCE 55.23 48.93 56.13 50.90
14 BCE 51.63 45.82 46.07 41.47
Table 4: Ablation study for the effect of visual grounding. We train the model separately on the ScanRefer and Multi3DRefer datasets.
Figure 3: The visualization results on ScanRefer. The green/red/blue colors indicate the correct/incorrect/ground truth boxes.
Figure 4: The visualization results on Scan2Cap. The input boxes are marked in blue.
Effectiveness of Position-Aware Video Representation. Table 3 presents the performance using different 3D position encoding (3D-PE) and coordinate aggregation strategies. The model was trained on the entire multi-task dataset, with 32 frames under the uniform frame sampling. We first assess the impact of different 3D-PE using average coordinate aggregation. The results indicate that the model can perform well without 3D-PE on 3D grounding and 3D QA tasks. This suggests that the video LLM paradigm can effectively understand spatial relationships across multiple frames. However, for dense captioning tasks, which require locating objects within video sequences based on query bounding boxes, the absence of 3D-PE results in significantly degraded performance. Additionally, MLP position encoding proves more effective for grounding tasks, while Sin PE works better for others. We then evaluate different coordinate aggregation strategies using Sin position encoding. Since the 3D coordinates are more complex and variable than those in 2D images, simply using the coordinates at the center of a 2D patch doesn’t accurately reflect the spatial location. As shown in the result, using average 3D coordinates better represents the 3D location within a patch, while using the minimum and maximum 3D coordinates within the patch is also a viable alternative.
[…]
Visualization. Figure 3 presents visualization results on the validation split of ScanRefer. The first and second rows indicate the rendered point clouds using uniform sampling and maximum coverage sampling with 16 frames. We observe that maximum coverage sampling usually provides more complete scenes than uniform sampling. For complex cases like (b-e), uniform sampling tends to miss smaller or peripheral objects, leading to prediction failures. Figure 4 presents visualization results on the Scan2Cap validation set. The first two rows show the rendered point clouds of the full video and the frames sampled using the MC strategy. With only 16 frames, nearly the entire scene’s information is captured. Meanwhile, using the proposed 3D-PE, the model accurately describes the specified target object.
5 Conclusion
In this paper, we propose a generalist model for 3D scene understanding, namely Video-3D LLM. It is built upon a Video LLM and incorporates 3D coordinates into video representations. By doing this, it aligns video representations with 3D physical world, thereby enabling a suite of 3D scene understanding tasks. In addition, we introduce a maximum coverage sampling strategy to balance computational costs with performance efficiency. Our extensive experimental results demonstrate that our model achieves state-of-the-art performance on several benchmarks.
Appendix
Appendix A Implementation Details
Dataset Statistics. We present the detailed statistics for training and testing data in Table 5 and 6, respectively. Following previous work [23, 22], we adopt the validation set for ScanRefer [6], Multi3DRefer [52], Scan2Cap [7], ScanQA [7], and the test set for SQA3D [36]. All data have been converted to LLaVA [34] format, and we conduct statistics in this format.
Evaluation Details. For ScanRefer [6], we select the object with the highest similarity as the prediction. For Multi3DRefer [52], we select objects with the highest probabilities such that their cumulative probability exceeds a given threshold p, which is empirically set to 0.25. For Scan2Cap [7], we follow [23] to evaluate the captioning performance by inserting “sos” and “eos” at the start and end of the prediction, respectively. Responses are generated using greedy sampling for 3D dense captioning and 3D question answering tasks.
Appendix B Detailed Comparison
SQA3D. We conduct a detailed evaluation on the test split of the SQA3D [36] dataset. As shown in Table 7, our model achieves the best performance on all categories of questions with an average EM at 58.86%, outperforming the previous state-of-the-art method LLaVA-3D [57] by 2.94% on the average EM.
Scan2Cap. As shown in Table 8, we provide a detailed comparison on the validation set of Scan2Cap [7]. During inference, we utilize the object proposals from [23], which include 50 predicted objects extracted with Mask3D [39] for each scan. From the table, we can see our method achieves the state-of-the-art on CIDEr and BLEU-4 at 83.77 and 42.43, respectively.
ScanRefer. We present a detailed comparison for ScanRefer [6] in Table 9. The table shows that our method reaches a peak of 58.12% Acc@0.25 and 51.72% Acc@0.5, surpassing ChatScene [22] by 2.6% and 1.5%, respectively.
Multi3DRefer. We follow previous work [52] to report the metrics across all question types, where “ZT” denotes zero-target, “ST” denotes single-target, “MT” denotes multi-target, “w/ D” and “w/o D” denote ‘with and without distractors, respectively. As shown in Table 10, our method outperforms previous methods on “ZT w/o D”, “ZT w/ D”, and “ST w/D” types. However, the performance for “MT” is lower than ChatScene [22], suggesting that our method still struggles to distinguish similar objects.
Data Count Scan Count Ques length Answer Length
ScanRefer [6] 36,665 562 24.9
Multi3DRefer [52] 43,838 562 34.8
Scan2Cap [7] 36,665 562 13.0 17.9
ScanQA [3] 26,515 562 13.7 2.4
SQA3D [36] 79,445 518 37.8 1.1
Table 5: Detailed statistics for training data. We report the average lengths for questions and answers, respectively.
Data Count Scan Count Ques length Answer Length
ScanRefer [6] (Val) 9,508 141 25.0
Multi3DRefer [52] (Val) 11,120 141 34.7
Scan2Cap [7] (Val) 2,068 141 13.0 18.7
ScanQA [3] (Val) 4,675 71 13.8 2.4
SQA3D [36] (Test) 3,519 67 36.3 1.1
Table 6: Detailed statistics for testing data. We report the average lengths for questions and answers, respectively.
Method Test set Avg.
What Is How Can Which Others
SQA3D [36] 31.6 63.8 46.0 69.5 43.9 45.3 46.6
3D-VisTA [58] 34.8 63.3 45.4 69.8 47.2 48.1 48.5
LLaVA-Video[53] 42.7 56.3 47.5 55.3 50.1 47.2 48.5
Scene-LLM [19] 40.9 69.1 45.0 70.8 47.2 52.3 54.2
LEO [23] 50.0
ChatScene [22] 45.4 67.0 52.0 69.5 49.9 55.0 54.6
LLaVA-3D [57] 55.6
Video-3D LLM (Uniform) 51.1 72.4 55.5 69.8 51.3 56.0 58.6
Video-3D LLM (MC) 50.0 70.7 57.9 69.8 50.1 55.8 57.7
Table 7: Performance comparison on the test set of SQA3D [36].
Method @0.5
C B-4 M R
Scan2Cap [7] 39.08 23.32 21.97 44.48
3DJCG [4] 49.48 31.03 24.22 50.80
D3Net [8] 62.64 35.68 25.72 53.90
3D-VisTA [58] 66.9 34.0 27.1 54.3
LL3DA [11] 65.19 36.79 25.97 55.06
LEO [23] 68.4 36.9 27.7 57.8
ChatScene [22] 77.19 36.34 28.01 58.12
LLaVA-3D [57] 79.21 41.12 30.21 63.41
Video-3D LLM (Uniform) 83.77 42.43 28.87 62.34
Video-3D LLM (MC) 80.00 40.18 28.49 61.68
Table 8: Performance comparison on the validation set of Scan2Cap [7]. C, B-4, M, R represent CIDEr, BLEU-4, Meteor, Rouge-L, respectively.
Method Venue Unique Multiple Overall
Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5
ScanRefer [6] ECCV20 76.33 53.51 32.73 21.11 41.19 27.40
MVT [24] CVPR22 77.67 66.45 31.92 25.26 40.80 33.26
3DVG-Transformer [54] ICCV21 81.93 60.64 39.30 28.42 47.57 34.67
ViL3DRel [9] NeurIPS22 81.58 68.62 40.30 30.71 47.94 37.73
3DJCG [4] CVPR22 83.47 64.34 41.39 30.82 49.56 37.33
D3Net [8] ECCV22 72.04 30.05 37.87
M3DRef-CLIP [52] ICCV23 85.3 77.2 43.8 36.8 51.9 44.7
3D-VisTA [58] ICCV23 81.6 75.1 43.7 39.1 50.6 45.8
3D-LLM (Flamingo) [21] NeurIPS23 21.2
3D-LLM (BLIP2-flant5) [21] NeurIPS23 30.3
Grounded 3D-LLM [14] ArXiv24 47.9 44.1
PQ3D [59] ECCV24 86.7 78.3 51.5 46.2 57.0 51.2
ChatScene [22] NeurIPS24 89.59 82.49 47.78 42.90 55.52 50.23
LLaVA-3D [57] ArXiv24 54.1 42.2
Video-3D LLM (Uniform) 87.97 78.32 50.93 45.32 58.12 51.72
Video-3D LLM (MC) 86.61 77.02 50.95 44.96 57.87 51.18
Table 9: Performance comparison on the validation set of ScanRefer [6]. “Unique” and “Multiple” depends on whether there are other objects of the same class as the target object.
Method ZT w/o D ZT w/ D ST w/o D ST w/ D MT ALL
F1 F1 F1@0.25 F1@0.5 F1@0.25 F1@0.5 F1@0.25 F1@0.5 F1@0.25 F1@0.5
M3DRef-CLIP [52] 81.8 39.4 53.5 47.8 34.6 30.6 43.6 37.9 42.8 38.4
D3Net [8] 81.6 32.5 38.6 23.3 35.0 32.2
3DJCG [4] 94.1 66.9 26.0 16.7 26.2 26.6
Grounded 3D-LLM [14] 45.2 40.6
PQ3D [59] 85.4 57.7 68.5 43.6 40.9 50.1
ChatScene [22] 90.3 62.6 82.9 75.9 49.1 44.5 45.7 41.1 57.1 52.4
Video-3D LLM (Uniform) 94.7 78.5 82.6 73.4 52.1 47.2 40.8 35.7 58.0 52.7
Video-3D LLM (MC) 94.1 76.7 81.2 72.6 52.7 47.4 40.6 35.3 57.9 52.4
Table 10: Performance comparison on the validation set of Multi3DRefer [52]. ZT: zero-target, ST: single-target, MT: multi-target, D: distractor.
Method Venue EM B-1 B-2 B-3 B-4 ROUGE-L METEOR CIDEr
ScanQA [3] CVPR22 21.05 30.24 20.40 15.11 10.08 33.33 13.14 64.86
3D-VisTA [58] ICCV23 22.4 10.4 35.7 13.9 69.6
Oryx-34B [35] ArXiv24 38.0 24.6 37.3 15.0 72.3
LLaVA-Video-7B [53] ArXiv24 39.71 26.57 9.33 3.09 44.62 17.72 88.70
3D-LLM (Flamingo) [21] NeurIPS23 20.4 30.3 17.8 12.0 7.2 32.3 12.2 59.2
3D-LLM (BLIP2-flant5) [21] NeurIPS23 20.5 39.3 25.2 18.4 12.0 35.7 14.5 69.4
Chat-3D [45] ArXiv23 29.1 6.4 28.5 11.9 53.2
NaviLLM [55] CVPR24 23.0 12.5 38.4 15.4 75.9
LL3DA [11] CVPR24 13.53 37.31 15.88 76.79
Scene-LLM [19] ArXiv24 27.2 43.6 26.8 19.1 12.0 40.0 16.6 80.0
LEO [23] ICML24 11.5 39.3 16.2 80.0
Grounded 3D-LLM [14] ArXiv24 13.4 72.7
ChatScene [22] NeurIPS24 21.62 43.20 29.06 20.57 14.31 41.56 18.00 87.70
LLaVA-3D [57] arXiv24 27.0 14.5 50.1 20.7 91.7
Video-3D LLM (Uniform) 30.10 47.05 31.70 22.83 16.17 49.02 19.84 102.06
Video-3D LLM (MC) 29.50 46.23 31.22 22.71 16.28 48.19 19.36 100.54
Table 11: Performance comparison on the validation set of ScanQA [3]. EM indicates exact match accuracy, and B-1, B-2, B-3, B-4 denote BLEU-1, -2, -3, -4, respectively.
ScanQA. We test our model on the validation set of ScanQA [3]. Compared to previous top-tier models, our Video-3D LLM achieves a relative improvement of 10.7% and 11.9% on EM@1 and CIDEr, respectively.